### v1

# Addressing Noise and Efficiency Issues in Graph-Based Machine Learning Models From the Perspective of Adversarial Attack

thanks: Correspondence to wangyongyu1@jd.com
Addressing Noise and Efficiency Issues in Graph-Based Machine Learning Models From the Perspective of Adversarial Attack
Yongyu Wang🖂 JD Logistics
preprint: APS/123-QED
Given that no existing graph construction method can generate a perfect graph for a given dataset, graph-based algorithms are invariably affected by the plethora of redundant and erroneous edges present within the constructed graphs. In this paper, we propose treating these noisy edges as adversarial attack and use a spectral adversarial robustness evaluation method to diminish the impact of noisy edges on the performance of graph algorithms. Our method identifies those points that are less vulnerable to noisy edges and leverages only these robust points to perform graph-based algorithms. Our experiments with spectral clustering, one of the most representative and widely utilized graph algorithms, reveal that our methodology not only substantially elevates the precision of the algorithm but also greatly accelerates its computational efficiency by leveraging only a select number of robust data points.
I Introduction
In the domain of graph-based algorithms like spectral clustering, the initial phase entails the construction of a graph from the provided dataset. This graph is structured such that each node corresponds to an individual data point, while the edges delineate the interrelations among these points. Given the intrinsic uniqueness of each dataset, it is impractical to expect a universal graph construction method that can precisely cater to all datasets. Consequently, the graphs generated may contain a significant number of erroneous or superfluous edges, commonly referred to as noisy edges. These noisy edges can profoundly degrade the performance of graph algorithms.
Currently, the most widely used and universally acknowledged best method for graph construction is the k-nearest neighbor (k-NN) graph. In a k-NN graph, each node is connected to its k nearest neighbors. This method possesses a strong capability to capture the local manifold, which is why it has become the predominant graph construction technique for the majority of graph algorithms. However, within datasets, the distribution and characteristics of points are neither uniform nor consistent, rendering the use of a uniform k value for all points imprecise. Furthermore, the metric used to measure clustering between two points is also problematic. Whether it is Euclidean distance, cosine similarity, or any other distance metric, each has its limitations. Consequently, the distances measured are not always accurate. For instance, in recommendation systems, different metrics perform with varying degrees of effectiveness across different datasets. The relationship between the metric used and the task at hand is not clearly defined and often relies on trial and error. Therefore, the k nearest neighbors in a k-NN graph may not necessarily be the ones that should be connected.
Addressing this issue is extremely challenging, primarily because, for the task of capturing the underlying structure of a dataset using a graph, there is no ground truth solution. That is to say, there is no definitive answer as to whether an edge is redundant or erroneous. Thus, only heuristic methods are available to remove noisy edges from the graph. Among the most representative works in this area is the consensus method proposed by [1], which extracts consensus information from a given $k$-NN graph. In this method, edges with a consensus value below a certain threshold are pruned. However, this approach, while removing noise, also eliminates a substantial number of non-noisy, useful edges. [2] proposed a spectral framework to construct a sparse graph to preserve the spectrum of the original graph for spectral graph algorithms. However, the gains in algorithmic accuracy are still relatively modest.
This work introduces a novel method aimed at enhancing the noise resilience of graph algorithms. Unlike approaches that attempt to remove noisy edges from the graph, our method employs a spectral adversarial robustness evaluation method to identify a small amount of robust nodes that exhibit strong resistance to noise. We then utilize only these robust nodes to complete the graph analysis tasks. The majority of graph-based algorithms such a shave a time complexity of at least $O(n^{2})$, and many are $O(n^{3})$ [3], such as spectral clustering algorithms, where $n$ is the number of nodes in the graph. Our method can significantly reduce the eigen-decomposition cost in spectral clustering, the dominant computational burden of the algorithm. This reduction is achieved because, unlike traditional approaches that involve all nodes, our method only requires the robust nodes, thus substantially lowering the computational cost.
The main contributions of this work are as follows:
  1. 1.
    To our knowledge, we are the first to propose treating noisy edges in graphs as adversarial attacks and to adopt a spectral proactive defense approach to address this issue.
  2. 2.
    In contrast to existing graph topology optimization approaches that address the issue of noisy edges from the perspective of the edges themselves, our method approaches the problem from the node perspective. Recognizing the inherent challenge in discerning whether a specific edge in a graph is noise or necessary—an issue that is often undecidable—we propose a solution that focuses on identifying nodes that are not vulnerable to noisy edges. By doing so, we aim to reconstruct a robust dataset that is resilient to the presence of noisy edges.
  3. 3.
    We show that by utilizing only a small number of robust nodes, significant improvements can be achieved in both the accuracy and efficiency for graph-based algorithms.
II Preliminaries
II.1 Adversarial attack and defense
Adversarial attack aims to misled machine learning models by providing deceptive inputs, such as samples with intentional disturbances [4, 5], which are commonly known as adversarial examples. [6, 7, 8] have demonstrated that machine learning models are often highly vulnerable to adversarial attacks.
To defend against adversarial attacks, many methods have been proposed. These methods can be categorized into two types: reactive defenses and proactive defenses. Reactive defenses concentrate on detecting adversarial examples within the model’s inputs, as investigated by [9, 10, 11]. In contrast, proactive defenses seek to bolster the robustness of the models, making them less vulnerable to the influence of adversarial examples, with methods presented by [12, 13].
II.2 Spectral clustering
Spectral clustering is one of the most representative and widely applied graph-based algorithms. It can often outperform traditional clustering algorithms, such as k-means algorithms[14], due to its ability to extract structural features of the dataset from the graph representation. There are three common spectral clustering algorithms used in practice, i.e., unnormalized spectral clustering, normalized spectral clustering according to Shi and Malik [15], and normalized spectral clustering according [16]. These algorithms are quite similar, apart from using different graph Laplacians. Typical spectral clustering algorithms can be divided into three steps: 1) construct a similarity matrix according to the entire data set, 2) embed all data points into k-dimensional space using eigenvectors of k bottom nonzero eigenvalues of the graph Laplacian, and 3) perform k-means algorithm to partition the embedded data points into k clusters.
III Method
III.1 Examining Graph-Based Machine Learning Models from the Perspective of Adversarial Attacks
Machine learning models are fundamentally mechanisms that map inputs to outputs via feature transformation. For instance, deep neural networks distill the original feature vectors of data through successive layers. Similarly, algorithms like Support Vector Machines (SVM) and Support Vector Classification (SVC) utilize kernel functions to map the original feature space of data into a higher-dimensional feature space. Adversarial attacks on machine learning models aim to mislead this mapping process.
In graph-based machine learning algorithms, the graph plays a pivotal role as the algorithm extracts structural information from the graph to transform input data, thereby mapping the inputs to outputs. Therefore, from the perspective of adversarial attacks, perturbing the graph can disrupt the mapping process of the machine learning model.
Suppose there exists a ground-truth ’perfect’ graph for a graph algorithm. In that case, any discrepancies between the actual graph constructed by our graph-building algorithm and the ground-truth graph could be considered as adversarial attacks applied to the ground-truth graph. Take spectral clustering as an example. This algorithm relies on constructing a graph to capture the underlying data structure of the dataset to be clustered. Subsequently, it computes the graph’s Laplacian matrix and employs the eigenvectors of the Laplacian matrix for spectral embedding, which maps the original high-dimensional feature space data into a lower-dimensional spectral domain. For instance, with the MNIST dataset, where the original feature space is 784-dimensional, the bottom eigenvectors of the Laplacian matrix derived from the graph are used to map the data into a 10-dimensional space for clustering purposes.
Ideally, the constructed graph should accurately capture the dataset’s underlying structure, allowing the mapping process to yield perfectly accurate clustering results. However, in reality, this is unattainable; no graph construction algorithm can create a completely accurate graph. Therefore, for any graph produced by a graph construction algorithm, the numerous noisy edges it contains can be viewed as adversarial attacks against it. Therefore, adversarial defense methods can be employed to manage these noisy edges.
III.2 Proactive Defense Strategies from a Node Perspective to Mitigate the Impact of Noisy Edges
Given that the ideal connectivity of edges within a graph is perpetually unknown, previous research focused on managing noisy edges directly from the edge perspective has yielded limited success. At times, this approach has inadvertently resulted in the removal of essential edges, leading to a degradation in algorithmic accuracy. Consequently, in this paper, we propose addressing the issue of noisy edges from the perspective of nodes.
Graph-based machine learning models, much like other machine learning models, predominantly focus on tasks associated with data points, such as classifying these points or predicting values associated with them. Graph algorithms distinguish themselves by utilizing the relationships signified by edges to assist in accomplishing tasks that are centered on the nodes. Edges in a graph play a twofold role. While they contain significant structural information that can aid in achieving more accurate data transformations—such as those utilized by graph neural networks to harness the associative information between nodes for improved performance—they can also be detrimental. A multitude of incorrect or superfluous edges can indeed degrade the performance of machine learning models. By enhancing the nodes’ resilience to noisy edges, we can exploit the beneficial edges to boost algorithmic performance without being adversely affected by the harmful ones.
In this paper, we propose enhancing the nodes’ resilience to noisy edges as a means to achieve improved algorithm performance by utilizing beneficial edges while concurrently mitigating the negative impact of harmful edges, when both coexist within the graph.
III.3 Adversarial Robustness Evaluation
In order to identify those with strong resilience, we need to evaluate the robustness of each data point to noisy edges. [17] proposed that the adversarial robustness of a given machine learning model can be measured by examining the distortion between the manifolds of the input feature space and the output feature space, by leveraging the generalized Courant-Fischer theorem. In this section, we employ this method to evaluate the robustness of each data point in the spectral clustering model. The specific steps are as follows:
  • Given data set $D$ with $N$ samples $x_{1},...,x_{N}\in{R}^{d}$ and its number of clusters $k$, we first construct a $k$-NN graph $G_{\text{input}}$ to capture the data manifold in the original $d$-dimensional feature space.
  • We perform the spectral embedding step in the spectral clustering algorithm to map the data points from the original $d$-dimensional space into $k$-dimensional spectral space to obtain data set $U$ with the points in the embedded feature space.
  • We construct a $k$-NN graph $G_{\text{output}}$ to capture the data manifold in the embedded $k$-dimensional feature space.
  • Based on the generalized Courant-Fischer theorem, [17] has further shown that the generalized eigenpairs of $L^{+}_{output}L_{input}$ can be used to estimate the robustness of each point, where $L^{+}_{output}$ denotes the Moore–Penrose pseudoinverse of the graph Laplacian of $G_{\text{output}}$ and $L_{input}$ denotes the graph Laplacian matrix of $G_{\text{input}}$.To this end, we construct the following eigensubspace matrix $\mathbf{V}_{k}\in\mathbb{R}^{N\times k}$:
    $\mathbf{V}_{k}{=}\left[\mathbf{v}_{1}\sqrt{\lambda_{1}},\ldots,\mathbf{v}_{k}\sqrt{\lambda_{k}}\right],$ (1)
    where $\lambda_{1},\lambda_{2},\ldots,\lambda_{k}$ represent the first $k$ largest eigenvalues of $L^{+}_{output}L_{input}$ and $\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{k}$ are the corresponding eigenvectors.
  • Finally, a metric called spade score for evaluating the adversarial robustness of a specific node $i$ can be calculated as follows [17]:
    $\text{spade}(i)=\frac{1}{|\mathcal{N}_{(}i)|}\sum_{j\in\mathcal{N}_{(}i)}\|\mathbf{V}_{k}^{\top}\mathbf{e}_{i,j}\|_{2}^{2};$ (2)
    where $j\in\mathcal{N}_{(}i)$ denotes the $j$-th neighbor of node $i$ in graph $G_{\text{input}}$, and $\mathcal{N}_{(}i)\subseteq V$ denotes the node set including all the neighbors of node $i$. A larger $\text{spade}(i)$ implies that node $i$ is likely more vulnerable to adversarial attacks.
III.4 A Multi-Level Algorithm Framework Based on Robust Node Set
We calculate the robustness score for all data points and sort them in ascending order based on their robustness scores. We then select a small number of data points with the lowest spade score, which correspond to the highest robustness, to form a robust subset. After obtaining the robust node set, we perform spectral clustering exclusively on this set to partition it into the desired number of k clusters. For each cluster, we calculate its centroid. Subsequently, each non-robust data point not selected for the robust set is assigned to the cluster whose centroid is closest to the data point.
III.5 Complete Algorithm Flow and Complexity Analysis
IV Experiment
In this section, we apply the proposed method to standard $k$-NN graph and use unnormalized spectral clustering to demonstrate its effectiveness. We assess the efficacy of our proposed approach by evaluating its ability to minimize the adverse effects of noise on the quality of solutions, as well as its capacity to increase the operational efficiency of the algorithms. Experiments are performed using MATLAB running on Laptop.
IV.1 Data sets
Experiments are performed using the following three real-world benchmark data sets:
  • USPS: includes $9,298$ images of USPS hand written digits with $256$ attributes.
  • MNIST: the machine learning field’s most recognized benchmark, features 60,000 training and 10,000 test images of handwritten digits, each with 784 attributes. We evaluate our methods using its test set.
IV.2 Metrics
To assess the solution quality of spectral clustering, we use the accuracy metric. It is defined as:
$ACC=\frac{\sum\limits_{j=1}^{n}{\delta{(y_{i},map(c_{i}))}}}{{n}},$ (3)
where $n$ represents the total count of data instances within the dataset, $y_{i}$ denotes the ground-truth label as provided by the dataset, and $C_{i}$ signifies the label ascribed by the clustering algorithm. The function $\delta(x,y)$ is a delta function, stipulated as: $\delta(x,y)$=1 for $x=y$, and $\delta(x,y)$=0, otherwise. The function $map(\bullet)$ serves as a permutation mapping that correlates each cluster index $c_{i}$ with an equivalent ground truth label, a process which can be efficiently accomplished utilizing the Hungarian algorithm [18]. An elevated $ACC$ value is indicative of superior clustering performance.
IV.3 Compared Algorithms
We compare our method against both the baseline and the state-of-the-art techniques for handling noisy edges in graphs. The specifics are as follows:
  • $k$-nearest neighbor graph:We choose $k$ = 10;
  • Consensus method: the state-of-the-art technique for selecting neighborhoods to construct affinity graphs. This method strengthens the graph’s robustness by incorporating consensus information from various neighborhoods in a specified kNN graph to select strong neighborhoods [1];
  • Spectral edge sparsification method: The state-of-the-art method for detecting non-critical, misleading, and superfluous edges in the graph [2].
IV.4 Results Of Solution Quality
Table 1 shows the solution quality of graph-based spectral clustering algorithm on the USPS and the MNIST data sets.
Table 1: Clustering Accuracy (%)
Data Set $k$-NN Consensus Spectral Spar Ours
USPS 64.31 68.54 70.74 78.87
MNIST 59.68 61.09 60.09 70.40
The clustering outcomes of our approach are derived by selecting the top 2,000 and 1,500 nodes with the highest robustness from the USPS and MNIST data sets, respectively. It is evident that our method surpasses the baseline kNN algorithm by over 14% and 10% in accuracy on the USPS and MNIST data sets, respectively, demonstrating the effectiveness of our approach in mitigating the effects of noise. Furthermore, the exceptional results achieved by our approach, which surpasses the second-best denoising method by margins of 8% for the USPS and 9% for the MNIST data sets, validate the advantage of employing an algorithm on nodes with reduced sensitivity to noisy edges over existing methods that concentrate on the elimination of such edges.
Experimental results reveal that methods aimed at resolving noisy issues by removing noisy edges do manifest a clear improvement on the USPS data set, although they fall short of the enhancements our method provides when compared to the baseline. On the MNIST data set, however, their effects are marginal, with the consensus method and spectral sparsification achieving only 1.5% and 0.5% increases in accuracy, respectively. We conjecture that this is attributable to the MNIST data set containing a greater number of features than the USPS data set, which may include more non-robust features that can induce noisy edges. Consequently, the task of excising noisy edges from MNIST is significantly more formidable. Existing denoising methods face difficulties in purging noisy edges without also eliminating beneficial edges. In contrast, our approach, which involves selecting nodes that are robust to noise and executing graph algorithms exclusively among these robust nodes, serves as an attack-agnostic method. This renders it highly effective for both the USPS and MNIST data sets.
IV.5 Efficacy of Resolving Computational Bottleneck
In spectral clustering algorithms, the first step of constructing the kNN graph can be completed within a time complexity of $O(|n|log|n|)$ [19], where $n$ is number of nodes in the graph. The third step, $k$-means, can also be performed in linear time [20],. However, the second step, eigen-decomposition, has a time complexity of $O(n^{3})$, making it the computational bottleneck of the entire algorithm, as well as the dominant term in the time complexity analysis of the method.
In our method, eigen-decomposition is only performed for a subset composed of a small number of robust nodes, thereby significantly reducing the computational cost.
Figure 1 and Table 2 show the eigen-decomposition time of the original full data set and our selected robust node set. It can be observed that by utilizing the selected robust node set, eigendecomposition has been accelerated by a factor of 9 for USPS and 90 for MNIST, respectively. It is expected that the proposed method will be a key enabler for running computationally expensive graph algorithms in scenarios that require extremely fast response times and on devices with limited computational capabilities.
Table 2: Eigen-decomposition Time (s)
Data Set Original Node Set Robust Node Set
USPS 0.72 0.08
MNIST 6.35 0.07
Figure 1: Eigen-decomposition time of original data set and selected robust nodes .
V Conclusion
In this paper, we have introduced a perspective that views the noisy edges in graphs used by graph algorithms through the lens of adversarial attacks. Building upon this viewpoint, we proposed a method based on robust nodes to mitigate the impact of noisy edges, grounded in adversarial robustness evaluation. Experimental results demonstrate that our approach significantly outperforms state-of-the-art methods for handling noisy edges by a large margin.